OpenStudy (anonymous):

can someone please help me figure out the x and y intercepts of this quadratic equation? y=2(x+4)^2-3

OpenStudy (anonymous):

in the file)

OpenStudy (anonymous):

If i get it right, you want y=0 => 2(x+4)^2-3=0 => x=sqrt(3/2)-4 AND x=-sqrt(3/2)-4 and x=0 , so y=2*4^2-3 => y= 29 SO there are three points of intersect, A(sqrt(3/2)-4,0), B(-sqrt(3/2)-4,0) and C(0,29) Note sqrt is square root :-)

OpenStudy (anonymous):

@Vladislav I can't view that file because it says that i need to download something....and I can't :/ can you print screen the what was in the document so that I can see?

OpenStudy (anonymous):

ok)

OpenStudy (anonymous):

@Theo177 I understood how to find the y-axis but I'm still quite confused for the x-axis @Vladislav So that's the graph of the equation right?

OpenStudy (anonymous):

Yes. x-horizontal, у-vertical

OpenStudy (anonymous):

X axis has 2 intercect points. \[\sqrt{3/2} -4\] is just a number

OpenStudy (zzr0ck3r):

2(x+4)^2-3 = 2(x^2 +8x +16)-3 = 2x^2 + 16x +29 set = 0 2x^2 +16x + 29=0 devide the twos away x^2 +8x + 29/2 = 0 (x+4)^2 = -29/2 + 16 x+4)^2 = -29/2 + 32/2 (x+4)^2 = 3/2 (x+4) = +-sqrt(3/2) x = -4 + sqrt(3/2) and -4 - sqrt(3/2)

OpenStudy (zzr0ck3r):

-0r+ sqrt(3/2) -4 is two numbers

OpenStudy (anonymous):

@zzr0ck3r how did you get the number 16 in the 6th line?

OpenStudy (anonymous):

\(y\) intercept is what you get when you replace \(x\) by 0 i get \[y=2(0+4)^2-3=2\times 16-3=32-3=29\]

OpenStudy (anonymous):

y=2(x+4)^2-3 For x=0, replace x with 0

OpenStudy (anonymous):

\(x\) intercept you get by setting equal to zero and solve \[2(x+4)^2-3=0\] \[2(x+4)^2=3\] \[(x+4)^2=\frac{3}{2}\] \[x+4=\pm\sqrt{\frac{3}{2}}=\pm\frac{\sqrt{6}}{2}\] \[x=-4\pm\frac{\sqrt{6}}{2}\]

OpenStudy (anonymous):

Oh thank you @satellite73 for making that clear! but how did you get the square root of 3/2 into the square root of 6/2?